OpenStudy (ns36):

Find a power series representation for f(x) = ln(1-x) and find the radius of convergence. Please help!

OpenStudy (amistre64):

since f' = -1/(1-x) we can do long division to get that power series; then integrate it back up to ln(1-x)

OpenStudy (ns36):

Thanks but I still don't quite understand how you get the power series from the derivative of the function :(

OpenStudy (amistre64):

the long division creates a pwer series; an equivalent polynomial. integrateing that takes us back to ln(1-x)

OpenStudy (anonymous):

Remember the geometric series?\[ \frac{1}{1-x} = \sum_{n = 0}^{\infty} x^{n}\]for |x| < 1. So we get -(1/(1-x)), that is:\[- \sum_{n = 0}^{\infty} x^n = -1 -x -x^2...\]Integrate this to get ln(1-x)

OpenStudy (amistre64):

-1-x-x^2-x^3-x^4- .... -------------------- 1-x ) -1 (-1+x) ------ -x (-x+x^2) --------- -x^2 \[D[ln(1-x)]=\frac{-1}{1-x}\] \[D[ln(1-x)]=-(1+x+x^2+x^3+x^4+...)\] integrate both sides \[ln(1-x)=\int -(1+x+x^2+x^3+x^4+...) dx\] \[ln(1-x)=-(x+\frac{x^2}{2}+\frac{x^3}{3}+\frac{x^4}{4}+\frac{x^5}{5}+...)\]

OpenStudy (ns36):

Thanks! How does one find the radius of convergence for a power series?

OpenStudy (amistre64):

this gives us the summation of ln(1-x) as:\[\sum_{n=0}^{inf}\frac{-x^{n+1}}{n+1}\] the radius of convergence is the limit as n goes to inf of the ratio of an+1/an

OpenStudy (anonymous):

There is a theorem that states that the derivatives (and integration?) have the same radius of convergence also.

OpenStudy (amistre64):

\[lim\frac{a_{n+1}}{a_{n}}:\ \lim_{n\to inf}\frac{-x^{n+1}}{n+1}*\frac{n}{-x^{n}}\]

OpenStudy (amistre64):

\[lim\frac{-x*n}{n+1}\to\ |-x|\lim\frac{n}{n+1}\] \[|-x|*1\] set it up such that: \[|-x|<1\]and simplify such that |x|<R and since theres nothing to simplify: R = 1

OpenStudy (ns36):

Hahaha thanks so much for all of your help!

OpenStudy (ns36):

How would I plug in (x-.5) into the power series you found above to express ln2 as a sum of an infinite series?

OpenStudy (amistre64):

when x=-1; you get ln(2) right?

OpenStudy (amistre64):

i think when we go beyond the radius of convergence is when we get iffy in values